OpenStudy (anonymous):

Waitress Pay: Candice Colton is a banquet waitress. She is paid $263 per hour plus 15% of the total cost of the food and beverages she serves during the banquet. If, during a 5 hour shift, Candice earns $400, What was the total cost of the food and beverages she served?

OpenStudy (dmezzullo):

Ok so, 400 including the 263? or no

OpenStudy (dmezzullo):

@ikoreanx3

Directrix (directrix):

Is $263 per hour supposed to be $2.63 per hour? I ask because if Candice works 5 hours at $263 per hour, she will earn 263*5 = $1,315. Then, this --> If, during a 5 hour shift, Candice earns $400 -- would not make sense because Candice would have earned $1,315 before 15% of the food and beverage expenses were added.

OpenStudy (anonymous):

Yes, I think its 2.63

OpenStudy (dmezzullo):

I was confused about tht to

Directrix (directrix):

What I think might be the equation is this: $2.63*5 + .15(FB) = $400 where FB is the cost of the food and beverages. That would give the FB costs to be $2,579 which sounds high. I wonder if the $400 total is supposed to be $40. Any ideas?

OpenStudy (anonymous):

I wrote down the problem exactly from the etext. And, I double checked now.It is written correctly...the problem. o:

Directrix (directrix):

Then, I think this would be the equation: $2.63*5 + .15(FB) = $400 What do you think @dmezzullo

OpenStudy (dmezzullo):

Looks right to me @Directrix

OpenStudy (dmezzullo):

Tell us the answer and we will check it when u have it if u want @ikoreanx3

Directrix (directrix):

I posted this 5 or 6 "rectangles" up in the thread. That would give the FB costs to be **$2,579** which sounds high.

Directrix (directrix):

I don't know the options.

OpenStudy (anonymous):

Unfortunately its not a multiple choice question. so I do not know the answer... should I just pick another question..

OpenStudy (dmezzullo):

How did u get over 2,500 dollars @Directrix from the 263 an hr instead of 2.63?

OpenStudy (dmezzullo):

@ikoreanx3 No giving up we will get u through this question

OpenStudy (anonymous):

Thank you T T

OpenStudy (dmezzullo):

All of us here believe u can do it so No giving up.

OpenStudy (anonymous):

I tried to follow Direct's equation..

Directrix (directrix):

I used this: "Yes, I think its 2.63" @ikoreanx3 wrote that above.

Directrix (directrix):

@ikoreanx3 What are the options? Maybe we can work backwards.

Directrix (directrix):

Ok so, 400 including the 263? or no Here is an equation that ignores the 5 hours. 400 = 263 + .15x where x is the food and beverage costs That renders $913.33 approx as the food and beverage costs.

OpenStudy (anonymous):

you mean 2.63? right..?

OpenStudy (dmezzullo):

Yes^

OpenStudy (anonymous):

maybe that could be it

Directrix (directrix):

Could you post a picture of the problem?

OpenStudy (anonymous):

I wish. but my laptop is broken and just ordered a new one yesterday it should come by tomorrow.

OpenStudy (anonymous):

so im on my smartphone and I see no option to attach a file.

Directrix (directrix):

Look on your smartphone and see what the answer options are.

OpenStudy (dmezzullo):

haha^

Directrix (directrix):

If we knew the options, we could figure out what the problem is.

OpenStudy (anonymous):

there is no answer options..

Directrix (directrix):

Okay. We'll work the problem as stated: Waitress Pay: Candice Colton is a banquet waitress. She is paid $263 per hour plus 15% of the total cost of the food and beverages she serves during the banquet. If, during a 5 hour shift, Candice earns $400, What was the total cost of the food and beverages she served? 263*5 +.15x = 400 where x is the total cost of food and beverages served x = - 6100 dollars which cannot be correct.